# Nav-$E$: Navigation-Guided Early Exiting for Efficient Vision-Language Models in Autonomous Driving

### v1

@@ IV-D Traffic-light Decision Results
Table II reports the performance of traffic-light state recognition on the BOSCH dataset. Compared with their full-layer baselines, all models benefit substantially from the proposed Nav-$E$ strategy. For example, DS-Tiny improves from 71.29% to 95.80% accuracy while reducing latency from 0.098s to 0.084s. Similarly, DS-Small increases accuracy from 52.01% to 77.55% with latency reduced from 0.248s to 0.186s. The largest improvement is observed with LLaVA-7B, which rises from 52.10% to 76.93% accuracy while cutting inference latency from 0.030s to 0.025s.
These results demonstrate that decision-oriented tasks such as traffic-light recognition, which require semantic reasoning beyond low-level perception, can also benefit from navigation-guided early exits. By leveraging task priors, Nav-$E$ avoids over-computation in later layers while maintaining high reliability, confirming its practical potential for safety-critical decision modules in autonomous driving.
@@ IV-E Ablation on General Early-Exit Baseline
To verify the necessity of navigation-aware configurations, we conduct an ablation study by comparing Nav-$E$ against a general early-exit (general-$E$) baseline. The general-$E$ strategy is implemented by following percentage-based layer reductions that are commonly adopted in LLMs for handling complex tasks, where inference is terminated after a fixed fraction of layers. As shown in Table III, this baseline only provides marginal improvements over the full-layer inference, and in some cases even reduces accuracy. For example, pedestrian recognition improves slightly from 66.93% to 67.11%, while traffic-light accuracy decreases from 52.10% to 37.68%. In contrast, our Nav-$E$ consistently delivers significant gains across all tasks. By leveraging navigation priors to guide task-specific exit layers, Nav-$E$ raises pedestrian, vehicle, and animal recognition to nearly perfect accuracy (100.0%, 99.98%, and 100.0% respectively) with more than 50% latency reduction. For the decision-critical traffic-light recognition task, Nav-$E$ improves accuracy from 52.10% to 76.93% while reducing latency from 0.030s to 0.025s.
@@ IV-F Real-Vehicle Evaluation
To validate the practicality of Nav-$E$ in real-world deployment, we integrate LLaVA-7B into an autonomous driving stack running on Autoware.Universe and conduct on-vehicle experiments. Table IV compares the baseline full-layer inference with our Nav-$E$ configuration across three representative tasks: person recognition, traffic-light recognition, and vehicle recognition. The results show consistent latency reductions while maintaining task reliability. Specifically, for person detection, Nav-$E$ reduces the inference time from 0.609s (32 layers) to 0.361s by exiting at layer 18. For traffic-light recognition, latency decreases from 0.611s to 0.486s with an exit at layer 25, and for vehicle recognition, Nav-$E$ exits as early as layer 14, cutting latency from 0.608s to 0.295s.
These findings demonstrate that Nav-$E$ can be effectively deployed in real-vehicle settings, offering significant efficiency gains (up to nearly 50% reduction in latency) without compromising task performance. This confirms the feasibility of navigation-guided early exit as a practical mechanism for reducing on-board computation in autonomous driving systems.
@@ V Conclusion
In this work, we introduced Nav-$E$, a navigation-guided early-exit framework for Vision-Language Models in autonomous driving. By combining offline profiling with online context switching, Nav-$E$ leverages navigation priors to configure task-specific exit layers, adapting inference to upcoming driving scenarios in a task-aware manner. Extensive experiments on CODA, Waymo, and BOSCH show that Nav-$E$ reduces inference latency while improving recognition accuracy across perception and decision tasks. Furthermore, real-vehicle deployment with Autoware.Universe confirmed its practicality, nearly halving end-to-end latency compared to full-layer baselines. Overall, Nav-$E$ demonstrates how predictive navigation can be effectively coupled with early-exit sparsification, enabling resource-efficient and task-aware deployment of large multimodal models under real-time constraints. Beyond efficiency, Nav-$E$ also mitigates the problem of over-thinking in full inference, ensuring more stable and reliable predictions. Its design is lightweight and compatible with existing autonomous driving stacks, requiring no retraining of the underlying VLMs. Future work will extend this framework toward broader multimodal tasks and integrate it with additional model compression techniques to further enhance scalability.

### v2

@@ IV-D Traffic-light Decision Results
Table II reports the performance of traffic-light state recognition on the BOSCH dataset. Compared with their full-layer baselines, all models benefit substantially from the proposed Nav-EE strategy. For example, DS-Tiny improves from 71.29% to 95.80% accuracy while reducing latency from 0.098s to 0.084s. Similarly, DS-Small increases accuracy from 52.01% to 77.55% with latency reduced from 0.248s to 0.186s. The largest improvement is observed with LLaVA-7B, which rises from 52.10% to 76.93% accuracy while cutting inference latency from 0.030s to 0.025s.
These results demonstrate that decision-oriented tasks such as traffic-light recognition, which require semantic reasoning beyond low-level perception, can also benefit from navigation-guided early exits. By leveraging task priors, Nav-EE avoids over-computation in later layers while maintaining high reliability, confirming its practical potential for safety-critical decision modules in autonomous driving.
@@ IV-E Ablation on General Early-Exit Baseline
To verify the necessity of navigation-aware configurations, we conduct an ablation study by comparing Nav-EE against a general early-exit (general-EE) baseline. The general-EE strategy is implemented by following percentage-based layer reductions that are commonly adopted in LLMs for handling complex tasks, where inference is terminated after a fixed fraction of layers. As shown in Table III, this baseline only provides marginal improvements over the full-layer inference, and in some cases even reduces accuracy. For example, pedestrian recognition improves slightly from 66.93% to 67.11%, while traffic-light accuracy decreases from 52.10% to 37.68%. In contrast, our Nav-EE consistently delivers significant gains across all tasks. By leveraging navigation priors to guide task-specific exit layers, Nav-EE raises pedestrian, vehicle, and animal recognition to nearly perfect accuracy (100.0%, 99.98%, and 100.0% respectively) with more than 50% latency reduction. For the decision-critical traffic-light recognition task, Nav-EE improves accuracy from 52.10% to 76.93% while reducing latency from 0.030s to 0.025s.
@@ IV-F Real-Vehicle Evaluation
To validate the practicality of Nav-EE in real-world deployment, we integrate LLaVA-7B into an autonomous driving stack running on Autoware.Universe and conduct on-vehicle experiments. Table IV compares the baseline full-layer inference with our Nav-EE configuration across three representative tasks: person recognition, traffic-light recognition, and vehicle recognition. The results show consistent latency reductions while maintaining task reliability. Specifically, for person detection, Nav-EE reduces the inference time from 0.609s (32 layers) to 0.361s by exiting at layer 18. For traffic-light recognition, latency decreases from 0.611s to 0.486s with an exit at layer 25, and for vehicle recognition, Nav-EE exits as early as layer 14, cutting latency from 0.608s to 0.295s.
These findings demonstrate that Nav-EE can be effectively deployed in real-vehicle settings, offering significant efficiency gains (up to nearly 50% reduction in latency) without compromising task performance. This confirms the feasibility of navigation-guided early exit as a practical mechanism for reducing on-board computation in autonomous driving systems.
@@ V Conclusion
In this work, we introduced Nav-EE, a navigation-guided early-exit framework for Vision-Language Models in autonomous driving. By combining offline profiling with online context switching, Nav-EE leverages navigation priors to configure task-specific exit layers, adapting inference to upcoming driving scenarios in a task-aware manner. Extensive experiments on CODA, Waymo, and BOSCH show that Nav-EE reduces inference latency while improving recognition accuracy across perception and decision tasks. Furthermore, real-vehicle deployment with Autoware.Universe confirmed its practicality, nearly halving end-to-end latency compared to full-layer baselines. Overall, Nav-EE demonstrates how predictive navigation can be effectively coupled with early-exit sparsification, enabling resource-efficient and task-aware deployment of large multimodal models under real-time constraints. Beyond efficiency, Nav-EE also mitigates the problem of over-thinking in full inference, ensuring more stable and reliable predictions. Its design is lightweight and compatible with existing autonomous driving stacks, requiring no retraining of the underlying VLMs. Future work will extend this framework toward broader multimodal tasks and integrate it with additional model compression techniques to further enhance scalability.